How many mL of 10.8M HCl are required to make 150.0 mL of 3.00M acid?





1
Expert's answer
2021-02-11T04:53:31-0500

Let the volume to be determined be x

X × 10.8 = 150 ×3

X = 150 ×3 /10.8

X= 41.67mL